OpenStudy (anonymous):

solve each equation: 2^3y+1=sqrt2

OpenStudy (akshay_budhkar):

is it 3y+1 together?

OpenStudy (anonymous):

yes

OpenStudy (akshay_budhkar):

ok we can write sqrt 2 as 2^1/2 therefore we have 3y+1=1/2 i believe u can go forward with that can you?

OpenStudy (anonymous):

we lost the 2 at the beginning of the equation

OpenStudy (anonymous):

\[2^{3y+1}=\sqrt{2}\]

OpenStudy (akshay_budhkar):

yes

OpenStudy (akshay_budhkar):

so sqrt2 is also equal to 2^1/2

OpenStudy (anonymous):

ok

OpenStudy (akshay_budhkar):

um so u can go about it? or u still confused?

OpenStudy (anonymous):

do I solve for Y now

OpenStudy (akshay_budhkar):

yes

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

No I guess I am still confused